OpenStudy (anonymous):

how to prove each equation is not a trigonometric identity by producing a counterexample. sin theta cos theta = cot theta

hartnn (hartnn):

Just assume one value of theta, say, theta = 30 substitute in both sides of equation. if left side does not equal right side, then its not an identity.

OpenStudy (anonymous):

oh okay thank you. because to substitute both sides i put 2pie/3

OpenStudy (anonymous):

and then i got root 3/2 and -1/2 =1

OpenStudy (anonymous):

but im not sure what to do next

hartnn (hartnn):

if you are substituting x=30 degrees, it'll be x=pi/6 radians. find sin pi/6 and cos pi/6 then.

OpenStudy (anonymous):

how come pi/6 radians? is there a reason i can't use like pi/ 4 radians? just curious

hartnn (hartnn):

you can use any angle.... even pi/4 i just took a particular angle.

OpenStudy (anonymous):

oh i see. thank you again.

OpenStudy (anonymous):

but first you would start we degrees and then change it to radians right?

OpenStudy (anonymous):

*with

hartnn (hartnn):

if you know the ratios in degrees, then you can use degrees also. like sin =/ directly.

hartnn (hartnn):

sin 30 =1/2 *

OpenStudy (anonymous):

could you help me with problem too sec ^2 x - 1 = cos x/ csc x

hartnn (hartnn):

if you put x=pi/4 or 45 degrees, left side not equal to right side. so, not an identity.